# ReCap: Better Gaussian Relighting with Cross-Environment Captures

### v1

@@ 4.2 Comparison with Previous Methods
Scene-wise Comparison.
ReCap consistently outperforms other methods on all three metrics across various unseen environments, as shown in  Tab. 2, while maintaining comparable NVS performance. As further supported by the visual comparisons in Fig. 6 where ReCap demonstrates noticeable perceptual improvements over other methods, this indicates advantageous relighting quality and fidelity without sacrificing NVS performance.
A notable observation across all methods is the large performance gap between NVS using the learned environment map and relighting using the ground truth HDR map (bridge* vs bridge). This highlights the albedo-lighting ambiguity, where the learned lighting representation diverges from ground truth but still works effectively with the learned materials for NVS. Among all methods, 3DGS-DR [44] shows the smallest gap, nonetheless its performance on unseen environments remains limited. As shown by the second row in Fig. 6, this is because that 3DGS-DR tends to preserve a consistent tone across different environments, closely matching the training views and adapting only the specular reflections based on new lighting.
Object-wise Comparison.
While scene-wise comparison confirms our method’s ability to generalize across different lighting conditions, a more granular evaluation of how effectively it adapts to diverse object attributes is presented in Tab. 3. Across various test objects featuring diffuse and specular surfaces, ReCap shows the best overall performance. Especially with more specular objects (e.g., toaster, musclecar) where other methods find challenging, ReCap shows a unique advantage in decoupling rich reflected lighting details for accurate relighting.
Reconstructed Environments.
While specular objects pose significant challenges for NVS due to their complex view-dependent appearance[18, 44], they offer distinct advantages for relighting tasks. Specular surfaces act as natural light probes, providing abundant high-frequency, view-dependent details that facilitate more accurate environment map reconstruction. In Fig. 7, we display the reconstructed environment maps for objects with varying level of specularity. When the object is mostly diffuse, it’s challenging to disentangle lighting effects from the object’s intrinsic properties due to insufficient lighting cues.
@@ 4.3 Ablation Study
Shading function.
To demonstrate the effectiveness of our proposed shading function in Eq. 8, we compare the average relighting performance across unseen scenes in  Tab. 4. The new shading function yields the most substantial improvement by offering a more flexible material representation, with the two additional regularization terms also contributing positively to the final results. Additional visual examples are provided in the supplementary material.
@@
Extra training environments.
Up to this point, we have focused on a dual-environment setup. In this experiment, we examine how extra photometric supervision from additional training environments affects the relighting results. Due to the constraint of 200 training views per scene, we allocate 100 identical views per environment as we include more environments for training. To provide a basis for comparison with existing results, we additionally compare the use of identical views versus extra views.
Results in Tab. 5 highlight three key findings: 1) more views from additional environment yields better results than adding more views within the same environment; 2) using identical views across environments provides stronger feature decoupling; and 3) increasing the number of environments consistently improves relighting robustness, showing no performance plateau up to five training environments.
@@ 4.4 Limitation and discussion
While we achieve compelling relighting results, indirect illumination and subsurface scattering effects are not considered. As shown in Tab. 3, objects exhibiting strong shadows (e.g., hotdog, coffee) or translucent materials (e.g., ship) represent common challenging cases. Over-exposed regions also pose inherent challenges for opacity estimation to all Gaussian based relighting methods, where clipping highlights are incorrectly interpreted as transparent against white splatting backgrounds.
# Envs # Views Cam Poses PSNR SSIM LPIPS
1 100 identical 23.98 0.906 0.084
1 200 extra 24.07 0.908 0.081
2 200 identical 26.14 0.927 0.066
2 200 extra 25.52 0.930 0.069
3 300 identical 26.35 0.930 0.064
4 400 identical 26.68 0.931 0.062
5 500 identical 27.36 0.936 0.060

### v2

@@ 4.2 Comparison with Previous Methods
Single-env Setup. We compare single-env performance in the first group of Tab. 2. The substantial difference between NVS using the learned environment map and relighting using the ground truth HDR map (bridge* vs. bridge) highlights the albedo-lighting ambiguity: the learned lighting representation diverges from ground truth but still works effectively with the learned materials for NVS. While 3DGS-DR [46] has the smallest gap, its ability to generalize to unseen relighting conditions remains limited. As shown in the second column of Fig. 6, this is because it tends to reproduce training views, preserving the overall tone and adapting only the specular reflections to new lighting.
Dual-env Setup. The second group of Tab. 2 presents results from directly optimizing existing Gaussian models over dual-env inputs while maintaining a single learned lighting representation. This leads to mixed relighting performance and significant NVS artifacts since the learned environment is no longer well-defined. Therefore, we continue our analysis with their single-env results for the remainder of the discussion. In the third group, we compare against TensoIR [19] with their multi-light setup and further augment GShader [18] with the proposed multi-light querying. Since TensoIR [19] uses GT albedo scaling in their official code, we additionally report their scale-free performance for clearer comparison.
Qualitative Comparison. Fig. 6 compares NVS and relighting performance across single-env Gaussian relighting methods, two dual-env TensoIR [19] variants, and dual-env ReCap. Our ReCap achieves noticeable perceptual improvements over methods that do not rely on GT for relighting. TensoIR [19] produces accurate relighting with GT albedo scaling, but suffers from obvious tone-shift without it (e.g., chair) and struggles with metallic objects (e.g., toaster) due to their dielectric assumption.
Reconstructed Environments. While specular objects pose significant challenges for NVS due to their complex view-dependent appearance[18, 46], they offer distinct advantages for relighting tasks. Specular surfaces act as natural light probes, providing abundant high-frequency, view-dependent details that facilitate more accurate environment map reconstruction. In Fig. 7, we display the reconstructed environment maps for objects with varying level of specularity. For predominantly diffuse objects, disentangling lighting from intrinsic properties is challenging due to limited lighting cues. Notice how our learned light map for chair retains object color, highlighting this challenge.
@@ 4.3 Ablation Study
Shading Function. To demonstrate the effectiveness of our proposed shading function in Eq. 8, we compare the average relighting performance across unseen scenes in  Tab. 3. The new shading function yields the most substantial improvement by offering a more flexible material representation, with the two additional regularization terms also contributing positively to the final results. Additional visual examples are provided in the supplementary material.
@@
Extra training environments. So far, we have focused on dual-env setups. Here, we examine how extra photometric supervision from additional training environments affects the relighting. Due to the constraint of 200 training views per scene, we allocate 100 identical views per environment as we include more environments for training. To align with existing results, we also compare the use of identical views versus extra views. Results in Tab. 4 reveal three key findings: (i) Adding identical views from an extra environment is better than adding extra unique views within the same environment (row 2 vs. 3); (ii) Although identical views across environments theoretically provides stronger decoupling, the benefit of extra unique views is more pronounced in practice (row 3 vs. 4). It is likely that extra views already provide sufficient decoupling, especially with perfect pose calibration in synthetic datasets; and (iii) Increasing the number of environments consistently enhances relighting performance, with no observed plateau up to five.
# Envs # Views # Unique Views Cam Poses PSNR SSIM LPIPS
1 100 100 identical 23.98 0.906 0.084
1 200 200 extra 24.07 0.908 0.081
2 200 100 identical 26.14 0.927 0.066
2 200 200 extra 26.25 0.928 0.065
3 300 100 identical 26.35 0.930 0.064
4 400 100 identical 26.68 0.931 0.062
5 500 100 identical 27.36 0.936 0.060
Application on Real Captures. ReCap requires multiple sets of camera poses to align to a common coordinate. In Fig. 8, we demonstrate its practicality with two real-life examples from StanfordORB [24], which captures the same object in different environments and provides COLMAP-estimated poses for each environment. They also provide sparse cross-environment alignment pairs computed by SuperGlue [38]. We use these pairs to estimate the view transform matrix and account for scale ambiguity. While exact calibration is non-trivial, sufficiently accurate pose calibration proves to be feasible for the proposed application.
Figure 8: Relighting and NVS examples on StanfordORB: gnome, teapot.
@@ 4.4 Limitation and discussion
While we achieve compelling relighting results, indirect illumination and subsurface scattering effects are not considered. Over-exposed regions also present inherent challenges for opacity estimation in all Gaussian-based relighting methods, where clipped highlights are incorrectly interpreted as transparent against white splatting backgrounds. For easier pose calibration in practice, semi-controlled capturing, such as rotating the object within the same environment at a known angle, may be considered.
@@ 5 Conclusion
Acknowledgment. This research was supported by the Alexander von Humboldt Foundation.